OpenStudy (anonymous):

PLEASE help and explain if you can :) let g(x) = x^2+3 if x

OpenStudy (anonymous):

ok you need to check which condition -4 fits

OpenStudy (anonymous):

since \[-4\leq 3\] you know it fits the first condition, so you use the first rule

OpenStudy (anonymous):

so \[g(-4)=(-4)^2+3=16+3=19\]

OpenStudy (anonymous):

if you needed say \[g(6)\] you would check and see that \[5<6<8\] so it fits the third condition and therefore \[g(6)=6^2-1=35\]

OpenStudy (anonymous):

wait so is there multiple answers for this problem?

OpenStudy (anonymous):

oh wait no , its just 19! thank you!

OpenStudy (anonymous):

yes 19

OpenStudy (anonymous):

you take your input and see which condition on the right that it fits. that tells you which fromula to use